OpenStudy (anonymous):

sqrt 9 m^2 n^2 + 2 sqrt m^2 n^2 - 3mn sqrt mn 2mn

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

@inowalst

OpenStudy (solomonzelman):

please draw parenthesis to indicate where square roots start and stop.

OpenStudy (fibonaccichick666):

Agreed, it currently reads \[\sqrt {9 m^2 n^2 + 2 \sqrt {m^2 n^2 - 3mn} }\]

OpenStudy (fibonaccichick666):

if you add a \ {bracket} to the beginning and the end you can type it as you did.

OpenStudy (anonymous):



OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm \sqrt{9 m^2 n^2} + 2\sqrt{m^2 n^2} - 3mn }\)

OpenStudy (solomonzelman):

correct?

OpenStudy (solomonzelman):

first, \(\normalsize\color{royalblue}{ \rm \sqrt{9 m^2 n^2} + 2\sqrt{m^2 n^2} - 3mn }\) can be re-written as \(\normalsize\color{royalblue}{ \rm \sqrt{3^3 m^2 n^2} + 2\sqrt{m^2 n^2} - 3mn }\) \(\normalsize\color{royalblue}{ \rm \Downarrow }\) \(\normalsize\color{royalblue}{ \rm \sqrt{(3mn)^2} + 2\sqrt{m^2 n^2} - 3mn }\) (will assume positive m and n, to avoid the absolute value) \(\normalsize\color{royalblue}{ \rm \Downarrow }\) \(\normalsize\color{royalblue}{ \rm 3mn + 2\sqrt{m^2 n^2} - 3mn }\)

OpenStudy (solomonzelman):

what cancels out so far? (Hint: \(\normalsize\color{royalblue}{ \rm \color{red}{3mn }+ 2\sqrt{m^2 n^2} \color{red}{- 3mn} }\) )

OpenStudy (anonymous):

so its 2mn?

OpenStudy (fibonaccichick666):

you may or may not be correct, first tell us how you came to that conclusion.

OpenStudy (solomonzelman):

After you subtracted the 3mn 's , you remained with \(\normalsize\color{royalblue}{ \rm 2\sqrt{m^2n^2} }\)

OpenStudy (solomonzelman):

We are not saying you are right, or you are wrong, we just want to see your thinking that made you conclude to the answer (because, we hope you aren't just guessing from the answer choices)

OpenStudy (anonymous):

thanks